Its me again! In a picnic a group of students rented a hut FOR P600. TWO OTHER STUDENTS JOINED IN, AS A CONSEQUENCES, EACH OF THE OTHERS NEED TO PAY P10 LESS. HOW MANY WERE THERE IN THE ORIGINAL GROUP?
Answer by scott8148(5891)

let x="original number" __ 600/x=(600/(x+2))+10

multiplying by x(x+2) __ 600x+1200=600x+10x^2+20x

subtracting 600x+1200 and dividing by 10 __ 0=x^2+2x-120

factoring __ 0=(x+12)(x-10) __ 0=x+12 and 0=x-10